# Random generation of universal cycles and de Bruijn sequences

### v1

@@ Abstract.
random generation, universal cycle, de Bruijn sequence, weak order, subsets, permutations, orientable sequence
@@ 1. Introduction
Let $\mathbf{\Sigma}_{k}(n)$ denote the set of all strings of length $n$ over the alphabet $\{0,1,\ldots,k{-}1\}$. Let $\mathbf{S}$ denote a subset of $\mathbf{\Sigma}_{k}(n)$. The de Bruijn graph of $\bf S$}, denoted $G(\mathbf{S})$, is the directed graph where each vertex corresponds to a length-(n1n{-}1) prefix or suffix of a string in $\bf S$}, and there is a directed edge from vertex $u$ to vertex $v$ if $u=u_{1}u_{2}\cdots u_{n-1}$ and $v=u_{2}u_{2}\cdots u_{n}$. Each edge corresponds to a string $u_{1}u_{2}\cdots u_{n}$ in $\mathbf{S}$ and we label such an edge by $u_{n}$. For example, see Figure 1. Interestingly, the de Bruijn graph of $\bf S$} is identical to the Rauzy graph [21] of a similar set. Though, Rauzy graph are typically defined with respect to strings.
@@
A universal cycle for $\mathbf{S}$, is a cyclic string of length $|\mathbf{S}|$ that contains each string in $\mathbf{S}$ as a substring exactly once (including the wraparound); they exist if and only if $G(\mathbf{S})$ is Eulerian, that is, $G(\mathbf{S})$ contains an Euler cycle. For example, consider $\mathbf{S}=\{001,010,101,011,110,100\}$ and the de Bruijn graph $G(\mathbf{S})$ illustrated in Figure 1(b): the Euler cycle $00,01,11,10,01,10,00$ corresponds to the universal cycle $110100$ obtained by outputting the labels on the edges in the Euler cycle. In this paper, we are concerned with interesting subsets $\mathbf{S}$ whose underlying de Bruijn graph is Eulerian, i.e., $\mathbf{S}$ admits a universal cycle. In particular, we consider:
@@
The primary objective of this paper is to describe practical algorithms to generate universal cycles for these objects uniformly at random, while using exponential space to store an underlying de Bruijn graph. Applying the work in [17], we describe a generic Las Vegas algorithm that generates a random Euler cycle in any directed Eulerian graph by first generating a random arborescence (spanning in-tree). For each set $\mathbf{S}$, we generate a random string in $\mathbf{S}$ in order to seed the algorithm by selecting a root for the arborescence. The generic algorithm is presented in Section 3. Then, in Section 4 we consider each of the aforementioned sets $\mathbf{S}$ and provide (i) a discussion of how to generate a random element from $\mathbf{S}$ and (ii) experimental evidence for the average cover time to compute a random arborescence in $G(\mathbf{S})$.
Motivation.   This paper is motivated by a recent result from Lipták and Parmigiani [19] that generates random de Bruijn sequences, although not uniformly at random (see Section 4.3). In that paper, they compared their approach with an implementation of Fleury’s algorithm [12] to generate Euler cycles, modified by adding randomization. That implementation could not generate all possible de Bruijn sequences, however, it served “as the closest available method for comparison” [19]. Indeed, despite the vast literature on universal cycle constructions, and in particular, de Bruijn sequences, we also found no detailed discussion or resource regarding the generation of these sequences uniformly at random, other than a passing comment by Propp and Wilson in [20, p.172]. This paper attempts to fill this void, even though it is known that a random arborescence in a directed Eulerian graph can be used to generate a random Euler cycle  [17].
@@ 2. Preliminaries
A traversal that starts and ends with the same vertex, say $r$, and visits all the edges in $E$ generates an Euler cycle, i.e., the traversal does not “burn bridges” [12]. This means that the |V|1|V|-1 edges corresponding to the last edges on each adjacency list, except for $r$’s, form an arborescence (spanning in-tree) rooted at $r$. Using this well-known fact, all Euler cycles can be generated as follows:
@@ 3. Random Generation
In this section, the algorithm outlined in Section 2 to generate all possible Euler cycles in a directed graph is applied to generate a single universal cycle uniformly at random for a set $\mathbf{S}$ with an underlying Eulerian de Bruijn graph $G(\mathbf{S})$. The Las Vegas algorithm describe by Algorithm R below summarizes the approach from Kandel et al. [17], which extends the work from [4]. We note one difference in our presentation. The algorithm in [17], takes as input a (cyclic) sequence $\mathcal{S}$ that may contain duplicate length $n$ substrings. From this sequence, it constructs a de Bruijn graph that allows for multiple edges between vertices. The algorithm is then initialized by selecting a random rotation of $\mathcal{S}$, which effectively generates a random edge in the underlying graph. For our purposes, we do not have an initial sequence $\mathcal{S}$. Instead, for each set $\mathbf{S}$ considered in Section 4, we present an efficient algorithm to compute a random element in $\mathbf{S}$, which corresponds to a random edge in $G(\mathbf{S})$.
@@
For Eulerian graphs, an arborescence can be generated uniformly at random by performing a random backwards walk until every vertex is visited. The first time a vertex is visited, the edge is recorded as a tree edge in the random arborescence [17]. The running time of this step depends on the cover time of the random walk, which is the number of steps it takes to visit every vertex. Thus, Algorithm R requires exponential space, with respect to $n$ and $k$, and also requires exponential time before the first symbol can be output. The final step (4) of Algorithm R generates a random universal cycle for $\mathbf{S}$ in constant time per symbol.
The expected cover time of a directed graph is $\Omega\left(2^{|V|}\right)$ in the worst case [8]. Theorem 3 in [17] shows that the expected cover time of a directed Eulerian graph is at most $|V|^{2}|E|$. In the next section, however, we demonstrate that some de Bruijn graphs appear to have a significantly faster expected cover time. Though, we do not have a proof of this observation, nor have we found a proof in the literature.
There are algorithms to generate random arborescences that are not dependent on the expected cover time. Wilson [28] shows how to produce a random arborescence in time proportional to the maximum hitting time. The maximum hitting time is the maximum over all pairs of vertices $u$ and $v$ of the expected number of steps for a random walk to travel from $u$ to $v$. Propp and Wilson [20] show how to produce a random arborescence in time proportional to the smaller of the maximum hitting time and the mean hitting time. The mean hitting time is the average over all pairs of vertices $u$ and $v$ of the expected number of steps for a random walk to travel from $u$ to $v$.
@@ 4.1. Permutations, subsets, and multiset permutations
It is well-known that a random permutation can be generated by applying the Fisher-Yates shuffle [11]; an $O(n)$ time implementation is provided by Knuth [18, Algorithm P] based on a presentation by Durstenfeld [10]. As noted by Arndt [5], it is straightforward to apply the shuffle to generate an unbiased random multiset permutation $m_{1}m_{2}\cdots m_{n}$ in $\mathcal{O}(n)$ time as illustrated in Algorithm 1. Thus, $m_{1}m_{2}\cdots m_{n-1}$ is a random shorthand multiset permutation that can be used to seed Algorithm R for shorthand permutations, shorthand subsets, and more generally, shorthand multiset permutations.
@@
Universal cycles for shorthand permutations can be constructed in $\mathcal{O}(1)$-amortized time per symbol using $\mathcal{O}(n)$ space [16, 23]. An $\mathcal{O}(n)$-time successor rule is presented in [15], and an $\mathcal{O}(1)$-amortized time per symbol algorithm applying concatenation trees is presented in [24] that uses $\mathcal{O}(n^{2})$ space. Universal cycles for shorthand subsets can be constructed in $\mathcal{O}(1)$-amortized time per symbol using $\mathcal{O}(n)$ space [22]. Universal cycles for shorthand multiset permutations (strings with fixed content) can be constructed in $\mathcal{O}(1)$-amortized time per symbol using $\mathcal{O}(n)$ space [25].
@@ 4.2. Weak orders
and $W_{3}=13$. The number of weak orders where there is a $k$-way tie for first is given by ${n\choose k}W_{n-k}$ for $k<n$; there is 1 weak order when $k=n$. Thus, $W_{n}=\sum_{k=1}^{n}{n\choose k}W_{n-k}$. To generate a random weak order, we will apply this recurrence to group the strings of $\mathbf{W}(n)$ based on their content. Let $c_{i}$ denote the number of occurrences of the symbol $i$ in $\omega=w_{1}w_{2}\cdots w_{n}$. Since every weak order contains 1 as its smallest symbol and the largest symbol possible is $n$, we say that $(c_{1},c_{2},\ldots,c_{n})$ is the content (also known as the Parikh vector) of $\omega$. For example, if $n=3$ then each weak order has content $(1,1,1),(1,2,0),(2,0,1),$ or $(3,0,0)$. If we partition $\mathbf{W}(n)$ into subsets based on their content, we can order the subsets based on the lexicographic ordering of the corresponding content. The weak orders with exactly one $1$ comes first, followed by those with exactly two $1$s and so on. Thus, to generate a random weak ordering in $\mathbf{W}(n)$, we can (i) select a random integer $r$ in $1,W_{n}]$, (ii) apply the recurrence to determine the content $(c_{1},c_{2},\ldots,c_{n})$ based on the described ordering, and (iii) apply a Fisher-Yates shuffle to obtain a random multiset permutation (weak order) with content $(c_{1},c_{2},\ldots,c_{n})$ (see Section 4.1). Details are provided in Algorithm 2. Note that the integer $r$ does not uniquely determine the weak order being generated, however, it is possible to obtain this property by applying an unranking algorithm for multiset permutations instead of applying the shuffle.
@@
Universal cycles for weak orders can be constructed via a successor rule that generates the sequence in $\mathcal{O}(n)$ time per symbol using $\mathcal{O}(n)$ space [27]. By applying concatenation trees, they can be generated in $\mathcal{O}(1)$-amortized time using $\mathcal{O}(n^{2})$ space [24]. See the enumeration sequence A000670 for $W_{n}$ in the Online Encyclopedia of Integer Sequences [1].
@@
If an application does not require a sequence generated uniformly at random, an algorithm which applies a Burrows-Wheeler transform can be applied; it outputs each de Bruijn sequence with positive probability [19]. The algorithm also requires an exponential amount of time, with respect to $n$ and $k$, before the first symbol can be output and requires exponential space. It produces each symbol in $\mathcal{O}(\alpha(2^{n}))$-amortized time per symbol for $k=2$, where $\alpha(n)$ is the inverse Ackerman function. The first estimate on the mean discrepancy of de Bruijn sequences is obtained using this algorithm [19].
In Table 5, we compare the running times of the (non-uniform) C++ algorithm implemented by the authors of [19] and our (uniform) algorithm implemented in C; both implementations are available online for download at [2]. We average the running time over 10 iterations, not including the time it takes to output the sequence. [^1] Our experiments were run on an iMac desktop with an Apple M4 processor. Our experimental times are faster than those reported in [19], which is likely due to a faster processor. Interestingly, using the standard rand() function in the latter implementation resulted in a large cycle of generated bits, which did not allow all vertices to be visited during the “random” walk when $n\geq 28$. Thus, an alternate method for generating random bits had to be deployed.  [^2] We added a value corresponding to the number of times the current vertex had been visited to the output of rand(). If running time is a concern, the non-uniform result from [19] may be preferred for such applications. However, if a sequence is desired to be generated uniformly at random, the one presented in this paper should be considered. It is also important to note that the algorithm described in this paper is a “Las Vegas” algorithm which means that no upper bound on the running time can be given, as it is not guaranteed to terminate.
@@
Table 5. Comparing the average clock time in seconds over 10 iterations between the algorithm from [19] which generates a de Bruijn sequence non-uniformly at random, and Algorithm R which generates a de Bruijn sequence uniformly at random.
@@ 4.4. Weight-range de Bruijn sequences
Weight-range de Bruijn sequences can be constructed via an $\mathcal{O}(n)$ time per symbol successor rule when the minimum weight is 0, or the maximum weight is $(k-1)^{n}$ [GSWW20]. In the binary case, they can be constructed for any weight range in $\mathcal{O}(1)$-amortized time [SWW13]. When $k=2$ and $\ell+1=u$, weight-range de Bruijn sequences correspond to the universal cycles for (shorthand) subsets discussed in Section 4.1.
@@
A necklace class is an equivalence class of strings under rotation; we call the lexicographically smallest string in the class a necklace. The necklace class containing $\alpha$ is denoted $\alpha]$. For example, if $\alpha=0001$ then $\alpha]=\{0001,0010,0100,1000\}$. Let $\mathbf{N}_{k}(n,z)$ denote the set of all necklaces in $\mathbf{\Sigma}_{k}(n)$ with no $0^{z}$ substring for $z>1$. All such necklaces end with 1 when $z\leq n$. Let ${\mathbf{Z}_{k}(n,z)=\bigcup_{\alpha\in\mathbf{N}_{k}(n,z)}[\alpha]}$. It is known that ${\mathbf{Z}_{k}(n,z)}$ admits a maximal length universal cycle that does not contain the substring $0^{z}$ [7]. We call a maximum length universal cycle that does not contain $0^{z}$ as a substring, a de Bruijn sequence with forbidden $0^{z}$.
@@
The lexicographically smallest de Bruijn sequences with forbidden $0^{z}$ can be generated via a simple greedy algorithm [26]; it can also be generated efficiently by concatenating the aperiodic prefixes of necklaces with no $0^{z}$ substring as they appear in lexicographic order [26, 13]. An exponential number of such sequences can be efficiently generated by applying cycle-joining as described in [7].

### v3

# Las Vegas algorithms to generate universal cycles and de Bruijn sequences uniformly at random

@@ Abstract.
Las Vegas algorithm, universal cycle, de Bruijn sequence, weak order, subsets, permutations, orientable sequence
@@ 1. Introduction
Let $\mathbf{\Sigma}_{k}(n)$ denote the set of all strings of length $n$ over the alphabet $\{0,1,\ldots,k{-}1\}$. Let $\mathbf{S}$ denote a subset of $\mathbf{\Sigma}_{k}(n)$. The de Bruijn graph of $\bf S$}, denoted $G(\mathbf{S})$, is the directed graph where each vertex corresponds to a length-(n1n{-}1) prefix or a length-(n1n{-}1) suffix of a string in $\bf S$}; for each string $u_{1}u_{2}\cdots u_{n}$ in $\mathbf{S}$ there is a directed edge labeled $u_{n}$ from vertex $u=u_{1}u_{2}\cdots u_{n-1}$ to vertex $v=u_{2}u_{2}\cdots u_{n}$. For example, see Figure 1.
@@
A universal cycle for $\mathbf{S}$, is a cyclic string of length $|\mathbf{S}|$ that contains each string in $\mathbf{S}$ as a substring exactly once (including the wraparound); they exist if and only if $G(\mathbf{S})$ is Eulerian, that is, $G(\mathbf{S})$ contains an Euler cycle. For example, consider $\mathbf{S}=\{001,010,101,011,110,100\}$ and the de Bruijn graph $G(\mathbf{S})$ illustrated in Figure 1(b): the Euler cycle $00,01,11,10,01,10,00$ corresponds to the universal cycle $110100$ obtained by outputting the labels on the edges in the Euler cycle. In this paper we are concerned with interesting subsets $\mathbf{S}$ whose underlying de Bruijn graph is Eulerian, i.e., $\mathbf{S}$ admits a universal cycle. In particular, we consider:
@@
The primary objective of this paper is to describe practical algorithms to generate universal cycles for these objects uniformly at random, while using exponential space to store an underlying de Bruijn graph. We recall a generic Las Vegas algorithm from [KMUW96] that generates a random Euler cycle in any directed Eulerian graph by first generating a random arborescence (spanning in-tree). For each set $\mathbf{S}$, we generate a random string in $\mathbf{S}$ in order to seed the algorithm by selecting a root for the arborescence. The generic algorithm is presented in Section 3. Then, in Section 4 we consider each of the aforementioned sets $\mathbf{S}$ and provide (i) a discussion of how to generate a random element from $\mathbf{S}$ and (ii) experimental evidence for the average cover time to compute a random arborescence in $G(\mathbf{S})$.
Motivation.   This paper is motivated by a recent result from Lipták and Parmigiani [LP24] that generates random de Bruijn sequences, although not uniformly at random (see Section 4.3). In that paper, they compared their approach with an implementation of Fleury’s algorithm [fleury] to generate Euler cycles, modified by adding randomization. That implementation could not generate all possible de Bruijn sequences, however, it served “as the closest available method for comparison” [LP24]. Indeed, despite the vast literature on universal cycle constructions, and in particular, de Bruijn sequences, we also found no detailed discussion or resource regarding the generation of these sequences uniformly at random, other than a passing comment by Propp and Wilson in [wilson2, p.172]. However, it is well known that a random arborescence in a directed Eulerian graph can be used to generate a random Euler cycle  [KMUW96].
@@ 2. Preliminaries
A traversal that starts and ends with the same vertex, say $r$, and visits all the edges in $E$ generates an Euler cycle, i.e., the traversal does not “burn bridges” [fleury]. This means that the |V|1|V|-1 edges corresponding to the last edges on each adjacency list, except for $r$’s, form an arborescence (spanning in-tree) rooted at $r$. Using this well-known fact, all Euler cycles can be generated as follows:
@@ 3. Random Generation
In this section the algorithm outlined in Section 2 to generate all possible Euler cycles in a directed graph is applied to generate a single universal cycle uniformly at random for a set $\mathbf{S}$ with an underlying Eulerian de Bruijn graph $G(\mathbf{S})$. The Las Vegas algorithm describe by Algorithm R below summarizes the approach from Kandel et al. [KMUW96], which extends the work from [alts]. We note one difference in our presentation. The algorithm in [KMUW96], takes as input a (cyclic) sequence $\mathcal{S}$ that may contain duplicate length $n$ substrings. From this sequence, it constructs a de Bruijn graph that allows for multiple edges between vertices. The algorithm is then initialized by selecting a random rotation of $\mathcal{S}$, which effectively generates a random edge in the underlying graph. For our purposes, we do not have an initial sequence $\mathcal{S}$. Instead, for each set $\mathbf{S}$ considered in Section 4, we present an efficient algorithm to compute a random element in $\mathbf{S}$, which corresponds to a random edge in $G(\mathbf{S})$.
@@
For Eulerian graphs, an arborescence can be generated uniformly at random by performing a random backwards walk until every vertex is visited. The first time a vertex is visited, the edge is recorded as a tree edge in the random arborescence [KMUW96]. The running time of this step depends on the cover time of the random walk, which is the number of steps it takes to visit every vertex. The algorithm requires $\Theta(|\mathbf{S}|)$ space to store the graph. The final step (4) of Algorithm R generates a random universal cycle for $\mathbf{S}$ in constant time per symbol.
Theorem 3 in [KMUW96] shows that the expected cover time of a directed Eulerian graph is at most $|V|^{2}|E|$. In the next section, however, our experiments seem to indicate that some de Bruijn graphs appear to have a significantly faster expected cover time. We do not have a proof of this observation, nor have we found a proof in the literature.
There are algorithms to generate random arborescences that are not dependent on the expected cover time. Wilson [wilson1] shows how to produce a random arborescence in time proportional to the maximum hitting time. The maximum hitting time is the maximum over all pairs of vertices $u$ and $v$ of the expected number of steps for a random walk to travel from $u$ to $v$. Propp and Wilson [wilson2] show how to produce a random arborescence in time proportional to the smaller of the maximum hitting time and the mean hitting time. The mean hitting time is the average over all pairs of vertices $u$ and $v$ of the expected number of steps for a random walk to travel from $u$ to $v$.
@@ 4.1. Permutations, subsets, and multiset permutations
It is well known that a random permutation can be generated by applying the Fisher-Yates shuffle [FisherYates]; an $O(n)$ time implementation is provided by Knuth [knuth2, Algorithm P] based on a presentation by Durstenfeld [durstenfeld]. As noted by Arndt [arndt], it is straightforward to apply the shuffle to generate an unbiased random multiset permutation $m_{1}m_{2}\cdots m_{n}$ in $\mathcal{O}(n)$ time as illustrated in Algorithm 1. Thus, $m_{1}m_{2}\cdots m_{n-1}$ is a random shorthand multiset permutation that can be used to seed Algorithm R for shorthand permutations, shorthand subsets, and more generally, shorthand multiset permutations.
@@
Universal cycles for shorthand permutations can be constructed in $\mathcal{O}(1)$ amortized time per symbol using $\mathcal{O}(n)$ space [shorthand2, shorthand]. An $\mathcal{O}(n)$-time successor rule is presented in [karyframework], and an $\mathcal{O}(1)$ amortized time per symbol algorithm applying concatenation trees is presented in [concattree] that uses $\mathcal{O}(n^{2})$ space. Universal cycles for shorthand subsets can be constructed in $\mathcal{O}(1)$ amortized time per symbol using $\mathcal{O}(n)$ space [fixed-weight-binary-strings]. Universal cycles for shorthand multiset permutations (strings with fixed content) can be constructed in $\mathcal{O}(1)$ amortized time per symbol using $\mathcal{O}(n)$ space [fixed-content].
@@ 4.2. Weak orders
and $W_{3}=13$. The number of weak orders where there is a $k$-way tie for first is given by ${n\choose k}W_{n-k}$ for $k<n$; there is 1 weak order when $k=n$. Thus, $W_{n}=\sum_{k=1}^{n}{n\choose k}W_{n-k}$.
To generate a random weak order, we will apply this recurrence to group the strings of $\mathbf{W}(n)$ based on their content. Let $c_{i}$ denote the number of occurrences of the symbol $i$ in $\omega=w_{1}w_{2}\cdots w_{n}$. Since every weak order contains 1 as its smallest symbol and the largest symbol possible is $n$, we say that $(c_{1},c_{2},\ldots,c_{n})$ is the content (also known as the Parikh vector) of $\omega$. For example, if $n=3$ then each weak order has content $(1,1,1),(1,2,0),(2,0,1),$ or $(3,0,0)$. If we partition $\mathbf{W}(n)$ into subsets based on their content, we can order the subsets based on the lexicographic ordering of the corresponding content. The weak orders with exactly one $1$ comes first, followed by those with exactly two $1$s and so on. Thus, to generate a random weak ordering in $\mathbf{W}(n)$, we can (i) select a random integer $r$ in $1,W_{n}]$, (ii) apply the recurrence to determine the content $(c_{1},c_{2},\ldots,c_{n})$ based on the described ordering, and (iii) apply a Fisher-Yates shuffle to obtain a random multiset permutation (weak order) with content $(c_{1},c_{2},\ldots,c_{n})$ (see Section 4.1). Details are provided in Algorithm 2. Note that the integer $r$ does not uniquely determine the weak order being generated, however, it is possible to obtain this property by applying an unranking algorithm for multiset permutations instead of applying the shuffle.
@@
Universal cycles for weak orders can be constructed via a successor rule that generates the sequence in $\mathcal{O}(n)$ time per symbol using $\mathcal{O}(n)$ space [weakorder]. By applying concatenation trees, they can be generated in $\mathcal{O}(1)$ amortized time using $\mathcal{O}(n^{2})$ space [concattree]. See the enumeration sequence A000670 for $W_{n}$ in the Online Encyclopedia of Integer Sequences [oeis1].
@@
If an application does not require a sequence generated uniformly at random, an algorithm which applies a Burrows-Wheeler transform can be applied; it outputs each de Bruijn sequence with positive probability [LP24]. The algorithm requires $\Theta(2^{n})$ space and produces each symbol in $\mathcal{O}(\alpha(2^{n}))$ amortized time per symbol for $k=2$, where $\alpha(n)$ is the inverse Ackerman function. It is important to note that $\alpha(n)$ grows extremely slowly, with $\alpha(n)\leq 5$ for any $n$ of practical value. The first estimate on the mean discrepancy of de Bruijn sequences is obtained using this algorithm [LP24].
In Table 5, we compare the running times of the (non-uniform) C++ algorithm implemented by the authors of [LP24] and our (uniform) algorithm implemented in C; both implementations are available online for download at [dbseq]. We average the running time over 10 iterations, not including the time it takes to output the sequence. [^1] Our experiments were run on an iMac desktop with an Apple M4 processor. Our experimental times are faster than those reported in [LP24], which is likely due to a faster processor. Interestingly, using the standard rand() function in the latter implementation resulted in a large cycle of generated bits, which did not allow all vertices to be visited during the “random” walk when $n\geq 28$. Thus, an alternate method for generating random bits had to be deployed.  [^2] We added a value corresponding to the number of times the current vertex had been visited to the output of rand(). If running time is a concern, the non-uniform result from [LP24] may be preferred for such applications. However, if a sequence is desired to be generated uniformly at random, the one presented in this paper should be considered. It is also important to note that the algorithm described in this paper is a “Las Vegas” algorithm which means that no upper bound on the running time can be given, as it is not guaranteed to terminate.
@@
Table 5. Comparing the average clock time in seconds over 10 iterations between the algorithm from [LP24] which generates a de Bruijn sequence non-uniformly at random, and Algorithm R which generates a de Bruijn sequence uniformly at random.
@@ 4.4. Weight-range de Bruijn sequences
Weight-range de Bruijn sequences can be constructed via an $\mathcal{O}(n)$ time per symbol successor rule when the minimum weight is 0, or the maximum weight is $(k-1)^{n}$ [GSWW20]. In the binary case, they can be constructed for any weight range in $\mathcal{O}(1)$ amortized time [SWW13]. When $k=2$ and $\ell+1=u$, weight-range de Bruijn sequences correspond to the universal cycles for (shorthand) subsets discussed in Section 4.1.
@@
A necklace class is an equivalence class of strings under rotation; we call the lexicographically smallest string in the class a necklace. The necklace class containing $\alpha$ is denoted $\alpha]$. For example, if $\alpha=0001$ then $\alpha]=\{0001,0010,0100,1000\}$. Let $\mathbf{N}_{k}(n,z)$ denote the set of all necklaces in $\mathbf{\Sigma}_{k}(n)$ with no $0^{z}$ substring for $z>1$. All such necklaces end with 1 when $z\leq n$. Let ${\mathbf{Z}_{k}(n,z)=\bigcup_{\alpha\in\mathbf{N}_{k}(n,z)}[\alpha]}$. It is known that ${\mathbf{Z}_{k}(n,z)}$ admits a maximal length universal cycle that does not contain the substring $0^{z}$ [RLL2]. We call a maximum length universal cycle that does not contain $0^{z}$ as a substring, a de Bruijn sequence with forbidden $0^{z}$.
@@
The lexicographically smallest de Bruijn sequences with forbidden $0^{z}$ can be generated via a simple greedy algorithm [generalize-classic-greedy]; it can also be generated efficiently by concatenating the aperiodic prefixes of necklaces with no $0^{z}$ substring as they appear in lexicographic order [generalize-classic-greedy, smaller_universal_cycles]. An exponential number of such sequences can be efficiently generated by applying cycle-joining as described in [RLL2].